OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

options!!!!

OpenStudy (anonymous):

y-(-1)=4(x+3) y+1=4x+12 y=4x+12-1 y=4x+11

OpenStudy (anonymous):

y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

y + 1 = 4(x + 3)